# Model Context Protocol (MCP): Landscape, Security Threats, and Future Research Directions

### v2

@@ 5.2.3. Sandbox Escape
Sandboxing isolates the execution environment of MCP tools, restricting their access to critical system resources and protecting the host system from potentially harmful operations. However, sandbox escape vulnerabilities arise when attackers exploit flaws in the sandbox implementation, enabling them to break out of the restricted environment and gain unauthorized access to the host system. Once outside the sandbox, attackers can execute arbitrary code, manipulate sensitive data, or escalate privileges, compromising the security and stability of the MCP ecosystem. Common attack vectors include exploiting weaknesses in system calls, improperly handled exceptions, and vulnerabilities in third-party libraries. For instance, a malicious MCP tool could exploit unpatched vulnerabilities in the underlying container runtime to bypass confinement and execute commands with elevated privileges. Similarly, side-channel attacks may allow attackers to extract sensitive data, undermining the intended isolation of the sandbox. Examining real-world sandbox escape scenarios in MCP environments can provide valuable insights for strengthening sandbox security and preventing future exploitation.
5.3. Security Risks in the Update Phase
The update phase involves managing server versions, modifying configurations, and adjusting access controls. This phase introduces three critical risks: post-update privilege persistence, re-deployment of vulnerable versions, and configuration drift.
@@ 5.3.1. Post-Update Privilege Persistence
Post-update privilege persistence occurs when outdated or revoked privileges remain active after an MCP server update, allowing previously authorized users or malicious actors to retain elevated privileges. This vulnerability arises when privilege modifications, such as API key revocations or permission changes, are not properly synchronized or invalidated following server updates. If these outdated privileges persist, attackers may exploit them to maintain unauthorized access to sensitive resources or perform malicious operations. For example, in API-driven environments like GitHub or AWS, privilege persistence has been observed when outdated OAuth tokens or IAM session tokens remain valid after privilege revocation. Similarly, in MCP ecosystems, if a revoked API key or modified role configuration is not promptly invalidated after an update, an attacker could continue invoking privileged actions, potentially compromising the integrity of the system. Enforcing strict privilege revocation policies, ensuring privilege changes propagate consistently across all server instances, and implementing automatic expiration for API keys and session tokens are essential to reducing the likelihood of privilege persistence. Comprehensive logging and auditing of privilege modifications further enhance visibility and help detect inconsistencies that could indicate privilege persistence.
5.3.2. Re-deployment of Vulnerable Versions
MCP servers, being open-source and maintained by individual developers or community contributors, lack a centralized platform for auditing and enforcing security updates. Users typically download MCP server packages from repositories like GitHub, npm, or PyPi and configure them locally, often without formal review processes. This decentralized model increases the risk of re-deploying vulnerable versions, either due to delayed updates, version rollbacks, or reliance on unverified package sources. When users update MCP servers, they may unintentionally roll back to older, vulnerable versions to address compatibility issues or maintain stability. Additionally, unofficial auto-installers, such as mcp-get and mcp-installer, which streamline server installation, may default to cached or outdated versions, exposing systems to previously patched vulnerabilities. Since these tools often prioritize ease of use over security, they may lack version verification or fail to notify users about critical updates. Because security patches in the MCP ecosystem rely on community-driven maintenance, delays between vulnerability disclosure and patch availability are common. Users who do not actively track updates or security advisories may unknowingly continue using vulnerable versions, creating opportunities for attackers to exploit known flaws. For example, an attacker could exploit an outdated MCP server to gain unauthorized access or manipulate server operations. From a research perspective, analyzing version management practices in MCP environments can identify potential gaps and highlight the need for automated vulnerability detection and mitigation. On the other hand, there is also a pressing need to establish an official package management system with a standardized packaging format for MCP servers and a centralized server registry to facilitate secure discovery and verification of available MCP servers.
@@ 5.3.3. Configuration Drift
Configuration drift occurs when unintended changes accumulate in the system configuration over time, deviating from the original security baseline. These deviations often arise due to manual adjustments, overlooked updates, or conflicting modifications made by different tools or users. In MCP environments, where servers are typically configured and maintained locally by end-users, such inconsistencies can introduce exploitable gaps and undermine the overall security posture. With the emergence of remote MCP server support, such as Cloudflare’s hosted MCP environments, configuration drift becomes an even more pressing concern. Unlike local MCP deployments, where configuration issues may only affect a single user’s environment, configuration drift in remote or cloud-based MCP servers can impact multiple users or organizations simultaneously. Misconfigurations in multi-tenant environments may expose sensitive data, lead to privilege escalation, or inadvertently grant malicious actors broader access than intended. Addressing this issue requires the implementation of automated configuration validation mechanisms and regular consistency checks to ensure that both local and remote MCP environments adhere to secure baseline configurations.
@@ 6.1. Implications
The rapid adoption of MCP is transforming the AI application ecosystem, introducing new opportunities and challenges that have significant implications for developers, users, MCP ecosystem maintainers, and the broader AI community.
For developers, MCP reduces the complexity of integrating external tools, enabling the creation of more versatile and capable AI agents that can perform complex, multi-step tasks. By providing a standardized interface for invoking tools, MCP shifts the focus from managing intricate integrations to enhancing agent logic and functionality. However, this increased efficiency comes with the responsibility to ensure that MCP implementations are secure, version-controlled, and aligned with best practices. Developers must remain vigilant about maintaining secure tool configurations and preventing potential misconfigurations that could expose systems to vulnerabilities.
For users, MCP enhances the experience by enabling seamless interactions between AI agents and external tools, automating workflows across platforms such as enterprise data management and IoT integration. It reduces the need for manual operations and improves efficiency in handling complex tasks. However, as MCP servers gain deeper access to sensitive data and critical operations, users must remain vigilant about the risks posed by unverified tools and misconfigured servers. Careless installation or untrusted sources may cause data leaks, unauthorized actions, or system instability.
For MCP ecosystem maintainers, the decentralized nature of MCP server development and distribution introduces a fragmented security landscape. MCP servers are often hosted on open-source platforms, where updates and patches are community-driven and may vary in quality and frequency. Without centralized oversight, inconsistencies in server configurations and outdated versions can introduce potential vulnerabilities. As the MCP ecosystem evolves to support remote hosting and multi-tenant environments, maintainers must remain attentive to potential risks associated with configuration drift, privilege persistence, and re-deployment of vulnerable versions.
For the broader AI community, MCP unlocks new possibilities by enhancing agentic workflows through cross-system coordination, dynamic tool invocation, and collaborative multi-agent systems. MCP’s ability to standardize interactions between agents and tools has the potential to accelerate AI adoption across industries, driving innovation in fields such as healthcare, finance, and enterprise automation. However, as MCP adoption grows, the AI community must address emerging ethical and operational concerns, such as ensuring fair and unbiased tool selection, safeguarding sensitive user data, and preventing potential misuse of AI capabilities. Balancing these considerations will be essential to ensuring that MCP’s benefits are widely distributed while maintaining accountability and trust within the AI ecosystem.
@@ 6.2. Challenges
Despite its potential, MCP’s adoption brings forth a range of challenges that need to be addressed to ensure its sustainable growth and responsible development:
Lack of centralized security oversight. Since MCP servers are managed by independent developers and contributors, there is no centralized platform to audit, enforce, or validate security standards. This decentralized model increases the likelihood of inconsistencies in security practices, making it difficult to ensure that all MCP servers adhere to secure development principles. Moreover, the absence of a unified package management system for MCP servers complicates the installation and maintenance process, increasing the likelihood of deploying outdated or misconfigured versions. The use of unofficial installation tools across different MCP clients further introduces variability in server deployment, making it harder to maintain consistent security standards.
Authentication and authorization gaps. MCP currently lacks a standardized framework for managing authentication and authorization across different clients and servers. Without a unified mechanism to verify identities and regulate access, it becomes difficult to enforce granular permissions, especially in multi-tenant environments where multiple users and agents may interact with the same MCP server. The absence of robust authentication protocols increases the risk of unauthorized tool invocation and exposes sensitive data to malicious actors. Moreover, inconsistencies in how different MCP clients handle user credentials further exacerbate these security challenges, making it difficult to maintain a consistent access control policy across deployments.
Insufficient debugging and monitoring mechanisms. MCP lacks comprehensive debugging and monitoring mechanisms, making it difficult for developers to diagnose errors, trace tool interactions, and assess system behavior during tool invocation. Since MCP clients and servers operate independently, inconsistencies in error handling and logging can obscure critical security events or operational failures. Without robust monitoring frameworks and standardized logging mechanisms, identifying anomalies, preventing system failures, and mitigating potential security incidents becomes challenging, hindering the development of more resilient MCP ecosystems.
Maintaining consistency in multi-step, cross-system workflows. MCP allows AI agents to execute multi-step workflows by invoking multiple tools across different systems through a unified interface. Ensuring consistent context across successive tool interactions is inherently difficult due to the distributed nature of these systems. Without effective state management and error recovery mechanisms, MCP risks propagating errors or losing intermediate results, leading to incomplete or inconsistent workflows. Additionally, dynamic coordination across diverse platforms can introduce delays and conflicts, further complicating the seamless execution of workflows within MCP environments.
Scalability challenges in multi-tenant environments. As MCP evolves to support remote server hosting and multi-tenant environments, maintaining consistent performance, security, and tenant isolation becomes increasingly complex. Without robust mechanisms for resource management and tenant-specific configuration policies, misconfigurations can lead to data leakage, performance issues, and privilege escalation. Ensuring scalability and isolation is critical for MCP’s reliability in enterprise deployments.
Challenges in embedding MCP in smart environments. Integrating MCP into smart environments, such as smart homes, industrial IoT systems, or enterprise automation platforms, introduces unique challenges related to real-time responsiveness, interoperability, and security. MCP servers in these environments must handle continuous streams of data from multiple sensors and devices while maintaining low-latency responses. Moreover, ensuring seamless interaction between AI agents and heterogeneous device ecosystems often requires custom adaptations, increasing development complexity. Compromised MCP servers in smart environments can lead to unauthorized control over critical systems, threatening both safety and data integrity.
6.3. Recommendations for MCP stakeholders
To safeguard the long-term success and security of MCP, all stakeholders, including MCP maintainers, developers, researchers, and end-users, should implement best practices and proactively address evolving challenges within the ecosystem.
Recommendations for MCP maintainers. MCP maintainers play a critical role in establishing security standards, improving version control, and ensuring ecosystem stability. To reduce the risk of security vulnerabilities, maintainers should establish a formal package management system that enforces strict version control and ensures that only verified updates are distributed to users. Additionally, introducing a centralized server registry would enable users to discover and validate MCP servers more securely, reducing the risk of interacting with malicious or misconfigured servers. To further enhance security, maintainers should promote the adoption of cryptographic signatures for verifying MCP packages and encourage periodic security audits to identify and mitigate vulnerabilities. Moreover, implementing a secure sandboxing framework can help prevent privilege escalation and protect host environments from malicious tool executions.
Recommendations for developers. Developers integrating MCP into AI applications should prioritize security and resilience by adhering to secure coding practices and maintaining thorough documentation. Enforcing version management policies can prevent rollbacks to vulnerable versions, while thorough testing ensures reliable MCP integrations before deployment. To mitigate configuration drift, developers should automate configuration management and adopt infrastructure-as-code (IaC) practices. Additionally, implementing robust tool name validation and disambiguation techniques can prevent conflicts that lead to unintended behavior. Leveraging runtime monitoring and logging helps track tool invocations, detect anomalies, and mitigate threats effectively.
Recommendations for researchers. Given the decentralized nature of MCP server deployment and the evolving threat landscape, researchers should focus on conducting systematic security analyses to uncover potential vulnerabilities in tool invocation, sandbox implementations, and privilege management. Exploring techniques to enhance sandbox security, mitigate privilege persistence, and prevent configuration drift can significantly strengthen MCP’s security posture. In addition, researchers should investigate more effective approaches for version control and package management in decentralized ecosystems to reduce the likelihood of re-deploying vulnerable versions. Researchers can help MCP maintainers and developers stay ahead of emerging threats by developing automated vulnerability detection methods and proposing secure update pipelines. Another critical area for research is the exploration of context-aware agent orchestration in multi-tool environments. As MCP increasingly supports multi-step, cross-system workflows, ensuring state consistency and preventing tool invocation conflicts becomes paramount. Researchers can explore techniques for dynamic state management, error recovery, and workflow validation to ensure seamless operation in complex environments.
Recommendations for end-users. End-users should remain vigilant about security risks and adopt practices to safeguard their environments. They should prioritize using verified MCP servers and avoid unofficial installers that may introduce vulnerabilities. Regularly updating MCP servers and monitoring configuration changes can prevent misconfigurations and reduce exposure to known exploits. Properly configuring access control policies helps prevent privilege escalation and unauthorized tool usage. For users relying on remote MCP servers, choosing providers that follow strict security standards can minimize risks in multi-tenant environments. Promoting user awareness and encouraging best practices will enhance overall security and resilience.

### v3

@@ 5.3.2. Sandbox Escape
MCP defines a communication framework between hosts and external tools. It focuses on structured data exchange and limited parameter validation, such as verifying types or value ranges. The protocol does not include mechanisms for runtime isolation or privilege control. As a result, the security boundaries of an MCP server depend entirely on its hosting environment. If the host already holds high system privileges, a poorly configured server may create conditions similar to a sandbox escape. This is particularly serious in MCP because servers often run inside AI‑integrated hosts that have broad access to files, networks, and user contexts. Once this boundary is breached, a malicious server can indirectly influence the agent’s reasoning process, alter shared data, or trigger privileged actions without the user’s awareness.
Different types of MCP servers face different risks. A database server that accepts unrestricted queries can allow unintended write or delete operations. A file-system server that lacks directory restrictions may enable access to sensitive areas such as configuration folders or system files. A network server without proper outbound filtering may let attackers send unauthorized requests or exfiltrate data. Other MCP servers, such as those connected to cloud APIs or local applications, may have similar but context-specific exposures. Real-world incidents demonstrate how such risks can emerge. Many MCP servers run inside host applications such as Cursor or Cline. These hosts already have permissions to read and write files, access networks, and interact with external services. If a host exposes an MCP server on localhost without authentication, other local processes, including a Chrome extension, can connect to it and issue privileged commands. In such cases, the attacker can act as an authorized user and bypass both the host sandbox and normal operating system protections. Several measures can help reduce these risks. The server should limit network bindings to trusted local clients and require authentication for every connection. File-system access should be restricted to a specific directory. Network permissions should allow only approved domains. All input parameters should be validated and sanitized before execution. In addition, containers or process-level isolation can help confine tool execution, and continuous monitoring can detect abnormal activity during runtime.
5.3.3. Tool Chaining Abuse
Tool chaining abuse occurs when multiple low-risk MCP tools are implicitly combined by the model to perform unintended high-impact operations. As shown in Figure 13, a single natural language request can trigger a sequence of legitimate tool calls that collectively lead to data exfiltration, uch as listing files, reading configuration data, extracting credentials, and exporting results to a public location. Because each step operates within the model’s authorized permissions, these activities often evade traditional access control or policy enforcement mechanisms. This attack is characterized by its implicit orchestration, where the model autonomously plans and chains approved tools without explicit malicious code or user intent. It exploits the semantic flexibility of natural language instructions and the composability of MCP tools, turning routine operations into multi-stage attack pipelines. The result is a highly stealthy and policy-compliant form of misuse that challenges conventional security monitoring and permission-based defenses.
Figure 13. Example of a tool chaining abuse attack. Multiple benign tools are combined by the model in response to a single natural language request. The AI application first lists directories and identifies a .env file, then reads it to extract database credentials, executes a query using those credentials to obtain sensitive user data, and finally writes the exported results to a public path.
Mitigation requires stronger execution constraints on cross-tool composition. Possible measures include prohibiting implicit tool chaining without explicit user approval, limiting which outputs can be piped to other tools (e.g., file contents to network tools), and enforcing runtime anomaly detection for suspicious chains such as “local file access $\rightarrow$ external network transmission”.
5.3.4. Unauthorized Access
Unauthorized access in the MCP ecosystem arises when attackers can interact with servers, tools, or communication channels beyond their intended permissions. Unlike prompt‐level threats that exploit the model’s language reasoning, unauthorized access typically abuses weaknesses in authentication, transport, or session management. As shown in Figure 14, the leakage of an SSE session_id can allow an attacker to reuse a valid session and issue remote commands without re‑authentication, leading to persistent system control. Similar risks occur when MCP components expose unprotected HTTP or WebSocket endpoints where sensitive functions, such as file or database operations, can be triggered directly by unauthorized clients. Another common source of unauthorized interaction stems from insecure credential or token handling. When environment variables or configuration files contain access tokens or API keys that are unintentionally exposed, adversaries can impersonate legitimate clients and perform privileged operations. In some deployments, authorization design flaws such as the confused deputy problem or token passthrough further amplify the issue: a trusted proxy may unknowingly perform actions on behalf of an attacker, or upstream credentials may be forwarded to downstream services without proper audience validation, breaking isolation boundaries. Session hijacking represents another manifestation of the same weakness, poorly protected or long‑lived session identifiers can be intercepted and reused to impersonate valid users, extending unauthorized control over the system. Building effective defenses requires mandatory endpoint authentication, secure credential storage, token audience enforcement, and the periodic rotation or invalidation of active sessions to restore strict trust boundaries between users, tools, and remote services.
Figure 14. Example of unauthorized access. By specifying transport=‘‘sse’’, the MCP server starts a Server-Sent Events (SSE) service on the default port 8000 and returns a unique session_id when the /sse route is accessed. If this identifier is exposed, an attacker can reuse it to send arbitrary commands through the /messages/?session_id=<id> route without authentication, enabling unauthorized command execution.
5.4. Security Flaws
5.4.1. Re-deployment of Vulnerable Versions
MCP servers, being open-source and maintained by individual developers or community contributors, lack a centralized platform for auditing and enforcing security updates. Users typically download MCP server packages from repositories like GitHub, npm, or PyPi and configure them locally, often without formal review processes. This decentralized model increases the risk of re-deploying vulnerable versions, either due to delayed updates, version rollbacks, or reliance on unverified package sources. When users update MCP servers, they may unintentionally roll back to older, vulnerable versions to address compatibility issues or maintain stability. Additionally, unofficial auto-installers, such as mcp-get and mcp-installer, which streamline server installation, may default to cached or outdated versions, exposing systems to previously patched vulnerabilities. Since these tools often prioritize ease of use over security, they may lack version verification or fail to notify users about critical updates. Because security patches in the MCP ecosystem rely on community-driven maintenance, delays between vulnerability disclosure and patch availability are common. Users who do not actively track updates or security advisories may unknowingly continue using vulnerable versions, creating opportunities for attackers to exploit known flaws.
Empirical evidence from the MCP ecosystem supports this observation. As shown in Table 4, the number of GitHub stars for unofficial auto-installers continues to increase, reflecting a growing user base and dependence on automated installation tools. However, the corresponding number of actively maintained MCP servers, including those hosted or updated through these installers, has remained largely unchanged over the same period. Except for the Smithery CLI, which shows regular updates, most auto-installers reference static MCP server packages that have not received version revisions in months. This asymmetry indicates that while more users are deploying MCP servers through automated tools, these tools frequently install outdated codebases lacking the latest security patches. As a result, the likelihood of re-deploying vulnerable versions increases proportionally with the popularity of auto-installers.
From a research perspective, analyzing version management practices in MCP environments can identify potential gaps and highlight the need for automated vulnerability detection and mitigation. On the other hand, there is also a pressing need to establish an official package management system with a standardized packaging format for MCP servers and a centralized server registry to facilitate secure discovery and verification of available MCP servers.
@@ 5.4.2. Post-Update Privilege Persistence
Privilege persistence refers to a condition in which outdated or revoked credentials remain valid after an MCP server update, enabling previously authorized users or malicious actors to retain elevated privileges. This vulnerability typically arises when privilege modifications, such as API key revocations, session token expirations, or role reassignments, are not properly synchronized or invalidated following a server upgrade. If such obsolete credentials persist, attackers can continue to access sensitive resources or perform privileged operations beyond the intended scope of the updated configuration. Similar behaviors have long been observed in conventional cloud and web systems, where revoked OAuth tokens or inactive IAM sessions remain temporarily valid due to asynchronous cache updates or delayed revocation. Robbins (Slack Security Team, 2022) demonstrated in 2022 that adversaries could obtain passwordless persistence and privilege escalation in Microsoft Azure Active Directory through the misuse of Certificate-Based Authentication. In that case, incomplete revocation of trusted credentials enabled a malicious certificate authority to impersonate global administrators indefinitely, illustrating how cached trust states and delayed revocation may sustain unauthorized privileges even after administrative changes. The MCP ecosystem inherits these underlying weaknesses and tends to amplify their effects. Many MCP servers maintain persistent network connections, such as SSE and WebSocket streams, while sharing credentials across different tools or hosts. When these servers are hot-reloaded or automatically updated, they may reload binaries without fully reinitializing credential data or terminating active sessions, which allows previously valid tokens to remain effective. As a result, a common web security problem that arises from delayed credential revocation can manifest with greater persistence and broader propagation in distributed MCP deployments.
To reduce this risk, MCP implementations should ensure immediate revocation of credentials during updates and synchronize privilege changes across all active instances. Automatic expiration and renewal mechanisms tied to server versions can further prevent old credentials from persisting. Integrating with centralized validation services that provide real-time checks of token status can strengthen overall consistency. Moreover, maintaining detailed logs and auditing privilege transitions can improve transparency and facilitate anomaly detection. From a design standpoint, extending the MCP specification with parameters such as session_expiry or explicit revocation_events could enhance lifecycle security and help prevent privilege persistence inherited from traditional web environments.
@@ 5.4.3. Configuration Drift
Configuration drift occurs when unintended or uncoordinated changes accumulate in a system’s configuration, causing it to deviate from the intended security baseline. Such deviations often result from manual adjustments, partial updates, or conflicting modifications introduced by different tools or users. In MCP environments, where server instances are frequently configured and maintained locally, configuration drift becomes a significant source of vulnerability. Because MCP servers may integrate multiple client tools and manage diverse access policies or API keys, even small inconsistencies can propagate across dependent components, leading to unpredictable security states. The emergence of remote or cloud-hosted MCP deployments, such as those provided by Cloudflare, heightens this concern. In contrast to local setups, where drift typically compromises only one user’s environment, configuration inconsistencies in multi-tenant or federated MCP servers can simultaneously affect many users or organizations. A misaligned access policy, outdated capability definition, or inconsistent plugin permission scope may inadvertently expose sensitive data or escalate privileges across tenants. These issues are amplified when servers rely on cached configurations or perform asynchronous synchronization between distributed nodes.
Mitigating configuration drift requires maintaining alignment between the deployed runtime state and a defined configuration baseline. Automated validation tools can periodically compare active settings with canonical templates and flag deviations before they propagate further. Version-controlled configuration definitions, immutable runtime manifests, and signed policy descriptors can provide verifiable reference points to detect unauthorized or accidental divergence. Additionally, implementing robust rollback procedures and continuous compliance auditing helps ensure that both local and remote MCP environments adhere to secure configuration standards. From a design perspective, incorporating configuration integrity verification into the MCP specification itself, for example through cryptographic checksums or schema validation hooks, would further reduce the risk of silent drift and strengthen overall operational resilience.
@@ 6.1. Implications
The rapid adoption of MCP is transforming the AI application ecosystem, introducing new opportunities and challenges that span the full MCP server lifecycle described in Figure 3, from creation to deployment, operation, and maintenance. These developments carry important implications for developers, users, ecosystem maintainers, and the broader AI community, each facing distinct responsibilities and risks across different lifecycle stages.
For developers, MCP simplifies tool integration during the creation and deployment stages (§ 3.3.2, § 3.3.3), enabling more efficient design of agentic workflows and complex multi-step reasoning tasks. By providing standardized capability declarations and invocation interfaces, developers can focus on functionality rather than integration friction. However, this openness also exposes new attack surfaces, such as namespace typosquatting (§ 5.1.1), tool poisoning (§ 5.1.4), and rug pulls (§ 5.1.5) during metadata definition or capability declaration. These threats can lead to privilege escalation, supply chain compromise, or execution of malicious payloads. Developers should therefore implement provenance verification, use version-controlled releases, and apply static analysis or digital signing before server registration to ensure integrity.
For users, the operation stage (§ 3.3.4) significantly enhances usability by enabling AI agents to orchestrate cross-platform workflows involving data services, enterprise systems, and IoT devices. This reduces manual effort and improves productivity. Nonetheless, users are directly exposed to runtime threats such as indirect prompt injection (§ 5.2.2), command injection (§ 5.1.7), and unauthorized access (§ 5.3.4). Interacting with unverified servers or insecure endpoints can lead to leaked credentials or unauthorized actions. To reduce exposure, users should prioritize trusted MCP collections, verify digital signatures at installation, and rely on sandboxed MCP hosts to confine tool execution during session management.
For MCP ecosystem maintainers, decentralization introduces heterogeneity across the deployment and maintenance phases (§ 3.3.3, § 3.3.5). Community-driven updates often vary in quality and cadence, leading to inconsistencies in patch management and version control. This fragmentation increases the likelihood of vulnerable versions (§ 5.4.1), privilege persistence (§ 5.4.2), and configuration drift (§ 5.4.3), which may be exploited to retain outdated permissions, misconfigure environments, or expose sensitive resources. Regular mechanisms such as automated version checks, integrity auditing, and mandatory configuration validation should be integrated into MCP registries and collections to ensure ecosystem-wide resilience.
For the broader AI community, MCP provides a foundation for interoperability that supports large-scale collaboration and reuse of AI capabilities across industries, enhancing cross-agent coordination and accelerating innovation. However, the expansion of server interactions across multiple lifecycle stages also amplifies ethical and safety concerns. Ensuring fairness in tool selection, defending against dataset leakage, and maintaining accountability in automated decision workflows become shared priorities. Continuous community participation in auditing, reporting, and lifecycle security monitoring is essential to preserve transparency, prevent misuse of AI capabilities, and ensure equitable access.
@@ 6.2. Challenges
Despite its potential, MCP adoption introduces a series of pressing challenges that must be resolved to ensure sustainable growth, operational reliability, and responsible development.
Lack of centralized security oversight. Since MCP servers are primarily managed by independent developers, there is no central authority to audit security baselines or enforce uniform compliance. This decentralization leads to inconsistent patching, irregular vulnerability management, and difficulty enforcing best practices. Malicious actors can exploit this fragmentation through namespace typosquatting (§ 5.1.1) or tool name conflicts (§ 5.1.2) during the metadata definition and capability declaration stages, resulting in the installation of malicious servers or privilege escalation. Moreover, the absence of an official package management mechanism leads to version inconsistency and unverified updates, heightening the risk of deploying vulnerable versions (§ 5.4.1) or misconfigured releases (§ 5.4.3).
Authentication and authorization gaps. MCP currently lacks a standardized framework for authentication and authorization across clients and servers. Without a unified mechanism to establish trust, identity spoofing and session hijacking can occur. As shown in Table 3, threats such as unauthorized access (§ 5.3.4) and privilege persistence (§ 5.4.2) can emerge during the operation and maintenance phases, allowing attackers to impersonate legitimate agents or retain administrative privileges. In multi-tenant environments, poorly implemented role-based access control further increases information exposure and inconsistent privilege enforcement across different MCP clients.
Insufficient debugging and monitoring mechanisms. The current MCP specification provides limited capabilities for logging, auditing, or runtime introspection. As a result, developers struggle to trace errors, investigate anomalous behavior, or detect silent failures during execution. Without unified telemetry and audit trails, incidents such as command injection (§ 5.1.7) and cross-server shadowing (§ 5.1.6) may remain undetected, enabling stealthy intrusion and lateral movement across connected environments. The lack of comprehensive monitoring frameworks also complicates response procedures, making it difficult to identify configuration drift or payload corruption until after exploitation has occurred.
Maintaining consistency in multi-step, cross-system workflows. MCP’s design supports dynamic, multi-step tool chaining, but ensuring consistent session state and reliable recovery is challenging in distributed environments. Malicious users may exploit these complex execution paths through tool chaining abuse (§ 5.3.3) or indirect prompt injection (§ 5.2.2) attacks, manipulating intermediate data or injecting hostile prompts between tools. Without robust state validation, integrity checks, or rollback mechanisms, small execution errors can propagate across multiple agents, resulting in workflow inconsistency or partial process execution.
Scalability challenges in multi-tenant environments. As MCP infrastructure evolves toward remote hosting and multi-tenant deployment, isolating tenants while preserving performance and privacy becomes more difficult. Unsynchronized updates or shared runtime environments may cause data leakage, resource contention, or mismanagement of privileges. Vulnerabilities such as sandbox escape (§ 5.3.2) and credential theft (§ 5.3.1) present notable threats during tool invocation within resource-shared clusters. To achieve scalability and security simultaneously, standardized tenant isolation models and runtime sandboxing must become integral to MCP server design.
Challenges in embedding MCP in smart environments. The integration of MCP into smart homes, industrial IoT systems, and enterprise automation introduces distinctive risks. These deployments often operate under real-time constraints and involve continuous communication between agents and heterogeneous devices. Compromised edge or on-premise MCP servers can lead to critical safety failures, privilege abuse, or command spoofing. Attacks such as installer spoofing (§ 5.2.1) or preference manipulation (§ 5.1.3) can exploit insecure installation pipelines or unsafe default configurations to deploy backdoored servers. Implementing secure firmware update channels, enforcing strict authentication for device-level access, and integrating MCP-aware intrusion detection are essential steps toward safer deployment in connected environments.
6.3. Recommendations across the MCP server lifecycle
The MCP can achieve secure, stable, and scalable adoption only if appropriate mechanisms are established at each stage of its lifecycle. Recommendations are therefore presented according to the four main lifecycle phases, textitcreation, deployment, operation, and maintenance, as defined in Figure 3. Each stage entails distinctive configuration tasks and corresponding opportunities for functional enhancement and risk mitigation.
6.3.1. Recommendations for Creation Phase (§ 3.3.2).
The creation phase is the starting point of an MCP server’s lifecycle. It defines how the server identifies itself, what functions it provides, and how securely those functions are implemented. Improvements made here have long-lasting effects on reliability and security. Three practical measures can strengthen this stage. Automate metadata generation and verification. Developers should not write metadata by hand. Instead, the server’s name, version, and supported protocols should be generated automatically during the build process. Each build should create a manifest file that includes these elements and a digital fingerprint of the source code. Before release, a verification step should confirm that this manifest matches the compiled artifact. This approach ensures that every published server can be traced to a trusted source and that outdated or modified copies are easy to detect. Validate capability declarations before release. A server’s capabilities describe its accessible tools and resources. Each capability should include a clear description of its scope, allowed operations, and required permission level. Developers should run a pre‑release validation process that checks whether all declared capabilities are correctly implemented and that none exceed their intended boundaries. For example, a capability that reads files should not include write access unless explicitly required. This validation helps prevent configuration errors and limits accidental privilege escalation during later operation. Separate implementation modules and enforce input checking. Each capability should have its own implementation module with limited access to shared resources. This design prevents faults in one module from affecting others. Within each module, the server should run input checks before executing any operation. These checks verify parameter counts, data formats, and expected value ranges. If inputs do not meet these conditions, the request should terminate safely instead of causing runtime errors. Clear separation between modules also makes it easier to test, update, or replace specific capabilities without changing the entire server.
6.3.2. Recommendations for Deployment Phase (§ 3.3.3).
The deployment phase introduces an MCP server into a live environment where client systems and external resources can interact with it. This stage determines whether the server operates safely and consistently across different hosts. Produce verified and reproducible build packages. Developers should prepare each release in a controlled build environment. The process should record dependency versions, configuration parameters, and source identifiers. Each package should include a checksum or digital signature that allows later verification of its origin. Before deployment, maintainers should confirm that the package metadata and version record match the server registry. This procedure guarantees that all running instances correspond to approved source code and that no altered or partially built versions enter production. Deploy servers within application sandboxes. Every MCP server should install and run inside a dedicated sandbox environment. The sandbox limits file access, network communication, and operating‑system instructions to a predefined scope. Server type determines the isolation rules. For instance, a server that connects to a database can operate through a restricted account. The sandbox allows only read queries, such as SELECT, when the server’s declared capability is to retrieve information. Modification commands like UPDATE or DELETE remain disabled. The sandbox can also intercept query inputs, apply escaping rules, and reject statements containing unapproved keywords to reduce the risk of SQL injection. Similarly, a server that interacts with the local file system should have access only to one designated directory created during installation. The sandbox enforces this directory boundary so that any attempt to read or write outside that folder fails immediately. Logs, configuration files, or temporary data are all stored within this enclosed workspace. This level of isolation prevents unnecessary exposure of system paths and confines possible damage to a controlled environment even if an error occurs.
6.3.3. Recommendations for Operation Phase (§ 3.3.4).
The operation phase is when the MCP server runs continuously and interacts with users, clients, and external resources. During this stage, the server interprets user intent, invokes tools, and manages sessions. Real‑time monitoring and control are essential to keep the system stable and secure. Monitor intent analysis and command execution. The MCP host or client should inspect the commands that users or language models generate before they reach the server. Each request should pass through an intent filter that checks whether the command matches declared capabilities and complies with safety rules. For example, an attempt to run a file deletion command against a read‑only tool should be rejected and logged before execution. Real‑time monitoring can also record how the intent is parsed and which capability is triggered. This visibility helps developers detect misuse patterns or unexpected behavior caused by ambiguous prompts. Maintain sandbox enforcement for dynamic operations. The sandbox introduced during deployment must remain active whenever the server handles live requests. Runtime policies should adapt to the server’s functional role. A database‑focused server should continue to use a restricted database account that only allows read queries. If a user or agent sends an update or delete request, the sandbox intercepts it and returns a permission error. For a server that manages files, the sandbox should keep all read and write operations within its dedicated workspace directory. Any attempt to access files outside this path should trigger an alert and terminate the operation safely. Maintaining active sandbox enforcement during execution ensures that even valid sessions cannot exceed defined resource boundaries. Implement consistent session management and adaptive logging. Each user or process connecting to the server should have its own authenticated session. The server should issue short‑lived tokens that expire automatically after inactivity. Session records need to track which capabilities are used, which resources are accessed, and what outputs are generated. Logging should occur in real time with minimal performance impact. For example, when a user triggers several tool invocations in a workflow, the log should record each tool name, execution time, and result summary. These logs allow administrators to identify errors, verify legitimate usage.
6.3.4. Recommendations for Maintenance Phase (§ 3.3.5).
The maintenance phase keeps the MCP server secure and reliable after deployment. Updates, configuration changes, and audits all happen during this stage. Apply structured version control and automated updates. Each server should maintain a clear version history that records all code and configuration changes. Developers should tag stable releases and store them in a controlled repository. When new patches are available, an automated update system should authenticate the package, verify compatibility, and apply changes without altering local configurations. If the update fails, the server should roll back automatically to the last stable version. This process minimizes downtime and prevents unverified patches from entering production environments. Manage configuration changes through validation and review. Administrators often need to update credentials, API keys, or resource addresses. Each configuration change should follow a controlled workflow. Before activation, the server should validate the new parameters to confirm that formats, ranges, and references are correct. For example, if a database host address changes, a test query should run automatically to confirm that the connection remains secure. Every modification should be logged together with its author, timestamp, and reason. A simple approval step, where another maintainer confirms the change, can further reduce misconfigurations that lead to service interruptions. Conduct continuous auditing and anomaly detection. The server should record key activities such as user authentication, tool invocation, and resource access. These logs should be stored in a secure and tamper‑evident format. Automated analysis can scan logs for signs of unusual activity, such as repeated failed logins or unexpected permission escalations. When anomalies occur, the system should send alerts to maintainers and isolate the affected component. Regular audits also help track long‑term behavior trends. For example, if a tool begins to take longer to execute common queries, maintainers can detect early signs of performance degradation or resource leaks before they cause larger failures.